If a 20 ft bottle is filled with 7 m of gallons for every 3 ft depth, how much gallon will it take to fill the brink of the bottle?
Answer by jndarrell(58)

Keep the units on the same side of the proportions ( either numerator or denominator) on both sides. IE ft/m=ft/m If you set it up m/ft=ft/m, you will not get the correct answer.
3ft/7m = 20ft/'x'm
Cross multiply
140=3x
Divide both sides by 3 to get x alone
x=46.667
REMINDER::
**Do not forget to report your units with your answer, 46.667 gallons
**Check your instructions for rounding rules
**Check your work by plugging in the value for x into the equation and work it through
:)